### v1

@@ C.2 Further Analysis on General Auditing Results
The value of $\tilde{\epsilon}$ for each target model is the gap between its corresponding dashed line and the baseline one in Figure LABEL:fig:eval:comparison. This allows us to compute values of $\tilde{\epsilon}$ reported in Table 4. It is also interesting to note that the maximum value of $\{c\!+\!\epsilon\}_{\textrm{lb}}$ typically occurs at low recall. Even if we do not use the same metric (precision-recall as opposed to TPR at low FPR) this is coherent with the findings from Carlini et al. (2022b). We can detect more privacy leakage (higher $\{c\!+\!\epsilon\}_{\textrm{lb}}$ which leads to higher $\tilde{\epsilon}$) when making a few confident guesses about membership rather than trying to maximize the number of members found (i.e., recall). Figure LABEL:fig:preds_comparison, decomposes the precision in terms of the number of true positives and the number of predictions for a direct mapping to the propositions 1 and 2. These are non-negligible values of privacy leakage, even through the true value is likely much higher.
C.3 PANORAMIA in cases of no privacy leakage
@@
As illustrated in Figure LABEL:fig:gpt2_privacy_recall, there exist instances in whch the $c_{\textrm{lb}}$ is as large or even larger than $\{c\!+\!\epsilon\}_{\textrm{lb}}$ (such as in GPT-2_E12). This implies that, under certain conditions, the baseline could outperform the MIA in PANORAMIA. We can see three possible reasons for this behavior. (1) The synthetic data might not be of high enough quality. If there is an evident distinction between real and synthetic samples, the baseline cannot be fooled reducing the power of our audit. (2) The MIA in PANORAMIA is not powerful enough. (3) The target model’s privacy leakage is indeed low. For GPT-2_E92 and GPT-2_E46 in Figure LABEL:fig:gpt2_privacy_recall, we can claim that given the target model’s leakage level, the synthetic data and the MIA can perform the audit to a non-negligible power. For GPT-2_12, we need to examine the possible reasons more closely. Focusing on generated data, in the ideal case we would like the synthetic data to behave just like real non-member data on Figure LABEL:subfig:lossdist-gpt-2E12. If this was the case, $\{c\!+\!\epsilon\}_{\textrm{lb}}$ would go up (since the MIA uses this loss to separate members and non-members). Nonetheless, predicting changes in $c_{\textrm{lb}}$ remains challenging, as analogous loss distributions do not guarantee indistinguishability between data points themselves (also, we cannot necessarily predict how the loss distribution changes under the helper model $h$ in Figure LABEL:subfig:lossdist-h). Hence, reason (1) which corresponds to low generative data quality, seems to be a factor in the under-performance of PANORAMIA. Moreover, in the case GPT-2_12, we observe that real members and non-members have very similar loss distributions. Hence, factor (2) also seems to be at play, and a stronger MIA that does not rely exclusively on the plot may increase auditing power. Finally, the helper model $h$ seems to have a more distinct loss distribution when compared to GPT-2_12. If the MIA had access to both target and helper models, we would expect $\{c\!+\!\epsilon\}_{\textrm{lb}}$ to at least match $c_{\textrm{lb}}$.
C.4 Failure of PANORAMIA in the case of tabular data
Our proposed method PANORAMIA can suffer some interesting failure cases that can be well illustrated with tabular data. Figure LABEL:fig:fail_privacy illustrates an interesting case of failure observed during one of our experiments with the generative model being trained on only part of the target model $f$ training set. First, by taking a closer look at the precision-recall curve of both the baseline (blue) and PANORAMIA ’s MIA for MLP_E100 on figure LABEL:fig:fail_precision, we notice that the precision-recall curve of the MIA is mostly above the baseline. However, the baseline has higher maximum precision values at extremely low recall. As a result, we can see in Figure LABEL:fig:fail_privacy that the maximum value for $(c+\varepsilon)_{lb}$ is lower than $c_{lb}$ leading to failure in measuring privacy leakage. This is due to our independent selection of recall thresholds, choosing heuristically the highest value.
@@ C.4 Failure of PANORAMIA in the case of tabular data
The high precision at low recall in the baseline illustrates the fact that some real data can easily be distinguished from the synthetic ones. Instead of poor quality synthetic data as exposed in C.3, we are dealing with real observation that are actually “looking to real”. Investigating those observations reveals that they have extreme or rare values in one of their features. Indeed, the Adult dataset contains some extreme values (e.g., in the capital_gain column), such extreme values are rarely reproduced by the generative model. This leads to the baseline accurately predicting such observations as real with high confidence that translates into a high $c_{lb}$. This behavior is the opposite of what we generally expect in MIA in which extreme values are usually more vulnerable to MIA Carlini et al. (2022b). This is a potential drawback of our auditing scheme, while the presence of extreme or rare values in the dataset usually allows for good auditing results it can have a negative impact on PANORAMIA’s success.
C.5 Privacy Auditing of Overfitted ML Models
@@
To verify if PANORAMIA will indeed attribute a higher value of $\tilde{\epsilon}$ to more overfitted models, we train our target models for varying numbers of training epochs. The final train and test accuracies are reported in Table 2. For GPT-2 (as a target model), we report the train and validation loss as a measure of overfitting in Figure 3. Figures LABEL:fig:gpt_losses and LABEL:fig:celeb_losses show how the gap between member data points (i.e., data used to train the target models) and non-member data points (both real as well as generated non-members) increases as the degree of overfitting increases, in terms of loss distributions. We study the distribution of losses since these are the features extracted from the target model $f$ or helper model $h$, to pass respectively to PANORAMIA and the baseline classifier. The fact that the loss distributions of member data become more separable from non-member data for more overfitted models is a sign that the corresponding target model could leak more information about its training data. We thus run PANORAMIA on each model, hereafter presenting the results obtained.
@@
Results. In Figure LABEL:fig:eval:comparison_precision, we observe that more training epochs (i.e., more overfitting) lead to better precision-recall trade-offs and higher maximum precision values. Our results are further confirmed by Figure LABEL:fig:preds_comparison with PANORAMIA being able to capture the number of member data points better than the baseline $b$.
@@
In Table 4, we further demonstrate that our audit output $\tilde{\epsilon}$ orders the target models in terms of privacy leakage: higher the degree of overfitting, more memorization and hence a higher $\tilde{\epsilon}$ returned by PANORAMIA. From our experiments, we consistently found that as the number of epochs increased, the value of $\tilde{\epsilon}$ also increased. For instance, in the WikiText-2 dataset, since we observe $\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E92}}>\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E46}}>\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E12}}$, we would have $\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E92}}-c_{\textrm{lb}}>\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E46}}-c_{\textrm{lb}}>\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E12}}-c_{\textrm{lb}}$, which leads to $\tilde{\epsilon}_{\text{GPT-2\_E92}}>\tilde{\epsilon}_{\text{GPT-2\_E46}}>\tilde{\epsilon}_{\text{GPT-2\_E12}}$. Our experiment is coherent with the intuition that more training epochs lead to more over-fitting, leading to more privacy leakage measured with a higher value of $\tilde{\epsilon}$.
C.6 Privacy Auditing of Differentially-Private ML Models
@@
Methodology. We evaluate the performance of PANORAMIA on differentially-private ResNet-18 and Wide-ResNet-16-4 models on the CIFAR10 dataset under different target privacy budgets ($\epsilon$) with $\delta=10^{-5}$ and the non-private ($\epsilon=\infty$) cases. The models are trained using the DP-SGD algorithm (Abadi et al., 2016) using Opacus (Yousefpour et al., 2021), which we tune for the highest train accuracy on learning rate $lr$, number of epochs $e$, batch size $bs$ and maximum $\ell_{2}$ clipping norm (C) for the largest final accuracy. The noise multiplier $\sigma$ is computed given $\epsilon$, number of epochs and batch size. Both PANORAMIA and O(1) (Steinke et al., 2023) audits privacy loss with pure $\epsilon$-DP analysis.
@@
Results. Tables 7 and 8 summarizes the auditing results of PANORAMIA on different DP models. For ResNet-18, we observe that at $\epsilon=1,2,4,6$ (more private models) PANORAMIA detects no privacy loss, whereas at higher values of $\epsilon=10,15,20$ (less private models) and $\epsilon=\infty$ (a non-private model) PANORAMIA detects an increasing level of privacy loss with $\tilde{\epsilon}_{\epsilon=10}<\tilde{\epsilon}_{\epsilon=15}<\tilde{\epsilon}_{\epsilon=20}<\tilde{\epsilon}_{\epsilon=\infty}$, suggesting a higher value of $\epsilon$ correspond to higher $\tilde{\epsilon}$. We observe a similar pattern with Wide-ResNet-16-4, in which no privacy loss is detected at $\epsilon=1,2$ and higher privacy loss is detected at $\epsilon=10,15,20,\infty$. We also compare the auditing performance of PANORAMIA with that of O(1) (Steinke et al., 2023), with the conclusion drawn by these two methods being comparable. For both ResNet-18 and Wide-ResNet-16-4, O(1) reports values close to 0 (almost a random guess between members and non-members) for $\epsilon<10$ DP models, and higher values for $\epsilon=10,15,20,\infty$ DP models. The results suggest that PANORAMIA is potentially an effective auditing tool for DP models that has comparable performance with O(1) and can generalize to different model structures.
@@
ResNet18 $\epsilon=\infty$ PANORAMIA RM;GN 2.508 3.6698 1.161 -
O (1) RM;RN - - - 1.565
ResNet18 $\epsilon=20$ PANORAMIA RM;GN 2.508 3.6331 1.125 -
@@
ResNet18 $\epsilon=15$ PANORAMIA RM;GN 2.508 3.5707 1.062 -
@@
ResNet18 $\epsilon=10$ PANORAMIA RM;GN 2.508 2.8 0.3 -
O (1) RM;RN - - - 0.14
ResNet18 $\epsilon=6$ PANORAMIA RM;GN 2.508 1.28 0 -
@@
ResNet18 $\epsilon=4$ PANORAMIA RM;GN 2.508 1.989 0 -
@@
ResNet18 $\epsilon=2$ PANORAMIA RM;GN 2.508 2.065 0 -
O (1) RM;RN - - - 0.08
ResNet18 $\epsilon=1$ PANORAMIA RM;GN 2.508 1.982 0 -
@@
WRN-16-4 $\epsilon=\infty$ PANORAMIA RM;GN 2.508 2.9565 0.448 -

### v2

@@ D.2 Further Analysis on General Privacy Measurement Results
The value of $\tilde{\epsilon}$ for each target model is the gap between its corresponding dashed line and the baseline one in Figure LABEL:fig:eval:comparison_privacy. This allows us to compute values of $\tilde{\epsilon}$ reported in Table 3. It is also interesting to note that the maximum value of $\{c\!+\!\epsilon\}_{\textrm{lb}}$ typically occurs at low recall. Even if we do not use the same metric (precision-recall as opposed to TPR at low FPR) this is coherent with the findings from Carlini et al. (2022b). We can detect more privacy leakage (higher $\{c\!+\!\epsilon\}_{\textrm{lb}}$ which leads to higher $\tilde{\epsilon}$) when making a few confident guesses about membership rather than trying to maximize the number of members found (i.e., recall). Figure LABEL:fig:preds_comparison, decomposes the precision in terms of the number of true positives and the number of predictions for a direct mapping to the propositions 1 and 2. These are non-negligible values of privacy leakage, even though the true value is likely much higher.
D.3 PANORAMIA performance with increasing Test and Train Size $\mathbf{m}$
PANORAMIA: Increasing Test Set Size: We compare PANORAMIA with the O(1) audit (Steinke et al., 2023) (in input space, without canaries), which uses a loss threshold MIA. As shown in LABEL:subfig:main-test-inc-resnet100 and LABEL:subfig:panoramia_vs_O1_m_dp_increase in the CIFAR10 image dataset case, both for non-private and Differentially Private (DP) models, PANORAMIA can leverage much more data for its measurement (up to the whole training set size for training and testing the MIA), while in our setting O(1) is limited by the size of the test set (for non-members). We vary test size from $m=500$, up to $35k$. The experiment is repeated $10$ times with resampling of the test set each time to produce the average. Unless otherwise stated, all results are produced at a $95\%$ confidence.
D.4 PANORAMIA with Increasing Train Dataset Size
We also conduct experiments to show the effect of increasing training dataset size on the performance of PANORAMIA (as well as the baseline). In the CIFAR10 image dataset case, Figure 15, we see PANORAMIA performance slowly increase as the train dataset size increases. It reaches its max performance at dataset sizes $4500-5000,10k,20k-30k$ and slowly begins to stabilize at the higher dataset sizes. We vary train size from $m=500$, up to $30k$. The experiment is repeated $5$ times with resampling of the train set each time to produce the average. Unless otherwise stated, all results are produced at a $95\%$ confidence.
Figure 15: Effect on PANORAMIA and baseline with increasing train set size for CIFAR10, ResNet101 target model (trained for 100 epochs) (with union bound correction).
D.5 PANORAMIA in cases of no privacy leakage
@@
As illustrated in Figure LABEL:fig:gpt2_privacy_recall and Table 3, there exist instances in which the $c_{\textrm{lb}}$ is as large as $\{c\!+\!\epsilon\}_{\textrm{lb}}$. For instance, for GPT-2_E37, the upper bound in the confidence interval measured for $c_{\textrm{lb}}$ is larger than the lower bound of the measurement for $\{c\!+\!\epsilon\}_{\textrm{lb}}$. Although the measured $\tilde{\epsilon}$ is positive on average, but the lower bound of the interval is negative. For such cases, we don’t claim any detection of privacy leakage. This implies that, under certain conditions, the baseline could outperform the MIA in PANORAMIA. We can see three possible reasons for this behavior. (1) The synthetic data might not be of high enough quality. If there is an evident distinction between real and synthetic samples, the baseline cannot be fooled reducing the power of our audit. (2) The MIA in PANORAMIA is not powerful enough. (3) The target model’s privacy leakage is indeed low. For GPT-2_E150 and GPT-2_E75 in Figure LABEL:fig:gpt2_privacy_recall, we can claim that given the target model’s leakage level, the synthetic data and the MIA can perform the audit to a non-negligible power. For GPT-2_37, we need to examine the possible reasons more closely. Focusing on generated data, in the ideal case we would like the synthetic data to behave just like real non-member data on Figure LABEL:subfig:lossdist-gpt-2E12. If this was the case, $\{c\!+\!\epsilon\}_{\textrm{lb}}$ would go up (since the MIA uses this loss to separate members and non-members). Nonetheless, predicting changes in $c_{\textrm{lb}}$ remains challenging, as analogous loss distributions do not guarantee indistinguishability between data points themselves (also, we cannot necessarily predict how the loss distribution changes under the helper model $h$ in Figure LABEL:subfig:lossdist-h). Hence, reason (1) which corresponds to low generative data quality, seems to be a factor in the under-performance of PANORAMIA. Moreover, in the case GPT-2_37, we observe that real members and non-members have very similar loss distributions. Hence, factor (2) also seems to be at play, and a stronger MIA that does not rely exclusively on the plot may increase auditing power. Finally, the helper model $h$ seems to have a more distinct loss distribution when compared to GPT-2_37 (see LABEL:subfig:lossdist-h).
D.6 PANORAMIA with Tabular data
PANORAMIA $\tilde{\varepsilon}$ O(1) $\varepsilon$
MLP_E10 0 $\pm$ 0.142 0.655 $\pm$ 0.014
MLP_E100 0.062 $\pm$ 0.156 0.648 $\pm$ 0.015
Table 6: Privacy measurement with PANORAMIA and O(1) for models trained on tabular Adult dataset.
As we can see in Table 6, our proposed method PANORAMIA cannot detect any meaningful privacy leakage in the case of a small MLP classifier trained on tabular data.
As shown in Figure LABEL:fig:privacy_recall_tab there is no significant difference between the values$(c+\varepsilon)_{lb}$ $c_{lb}$ returned by PANORAMIA. As a result we get $\tilde{\varepsilon}=0$
@@ D.6 PANORAMIA with Tabular data
The high precision at low recall in the baseline as can be seen in Figure LABEL:fig:precision_recall_tab illustrates the fact that some real data can easily be distinguished from the synthetic ones. Instead of poor quality synthetic data , we are dealing with real observation that are actually “looking to real”. Investigating those observations reveals that they have extreme or rare values in one of their features. Indeed, the Adult dataset contains some extreme values (e.g., in the capital_gain column), such extreme values are rarely reproduced by the generative model. This leads to the baseline accurately predicting such observations as real with high confidence that translates into a high $c_{lb}$. Moreover, the audit values maximum values $(c+\varepsilon)_{lb}$ and $c_{lb}$ being achieved for a relatively small number of guesses corresponding to extreme values, the audits results vary a lot depending on the sampling or not of such extreme values in the audit test set. We believe this is part of the reason why we observe an high variance in the audit results leading to no significant privacy leakage detected by PANORAMIA.
This is a potential drawback of our auditing scheme, while the presence of extreme or rare values in the dataset usually allows for good auditing results it can have a negative impact on PANORAMIA’s success.
D.7 Privacy Auditing of Overfitted ML Models
@@
To verify if PANORAMIA will indeed attribute a higher value of $\tilde{\epsilon}$ to more overfitted models, we train our target models for varying numbers of training epochs. The final train and test accuracies are reported in Table 1. For GPT-2 (as a target model), we report the train and validation loss as a measure of overfitting in Figure 13. Figures  LABEL:fig:celeb_losses and  LABEL:fig:gpt_losses show how the gap between member data points (i.e., data used to train the target models) and non-member data points (both real as well as generated non-members) increases as the degree of overfitting increases, in terms of loss distributions. We study the distribution of losses since these are the features extracted from the target model $f$ or helper model $h$, to pass respectively to PANORAMIA and the baseline classifier. The fact that the loss distributions of member data become more separable from non-member data for more overfitted models is a sign that the corresponding target model could leak more information about its training data. We thus run PANORAMIA on each model, hereafter presenting the results obtained.
@@
Results. In Figure LABEL:fig:eval:comparison_precision, we observe that more training epochs (i.e., more overfitting) lead to better precision-recall trade-offs and higher maximum precision values. Our results are further confirmed by Figure LABEL:fig:preds_comparison with PANORAMIA being able to capture the number of member data points (positive predictions) better than the baseline $b$.
@@
In Table 3, we further demonstrate that our audit output $\tilde{\epsilon}$ orders the target models in terms of privacy leakage: higher the degree of overfitting, more memorization and hence a higher $\tilde{\epsilon}$ returned by PANORAMIA. From our experiments, we consistently found that as the number of epochs increased, the value of $\tilde{\epsilon}$ also increased. For instance, in the WikiText dataset, since we observe $\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E150}}>\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E75}}>\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E37}}$, we would have $\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E150}}-c_{\textrm{lb}}>\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E75}}-c_{\textrm{lb}}>\{c\!+\!\epsilon\}_{\textrm{lb}}^{\text{GPT-2\_E37}}-c_{\textrm{lb}}$, which leads to $\tilde{\epsilon}_{\text{GPT-2\_E150}}>\tilde{\epsilon}_{\text{GPT-2\_E75}}>\tilde{\epsilon}_{\text{GPT-2\_E37}}$. Our experiment is coherent with the intuition that more training epochs lead to more over-fitting, leading to more privacy leakage measured with a higher value of $\tilde{\epsilon}$.
Figure 17: CIFAR-10, $\{c+\epsilon\}_{lb}$ when varying privacy leakage with increasing model parameters over one experiment run.
D.8 Models of varying complexity:
Carlini et al. (2021) have shown that larger models tend to have bigger privacy losses. To confirm this, we also audit ML models with varying numbers of parameters, from a $\approx 4M$ parameters Wide ResNet-28-2, to a $25.5M$ parameters ResNet-50, and a $44.5M$ parameters ResNet-101. Figure 17 shows that PANORAMIA does detect increasing privacy leakage, with $\tilde{\epsilon}_{wide-resnet}\leq\tilde{\epsilon}_{resnet50}\leq\tilde{\epsilon}_{resnet101}$.
D.9 Privacy Auditing of Differentially-Private ML Models
Methodology.
@@ D.9 Privacy Auditing of Differentially-Private ML Models
For the image data modality, we evaluate the performance of PANORAMIA on differentially-private ResNet-18 and Wide-ResNet-16-4 models on the CIFAR10 dataset under different target privacy budgets ($\epsilon$) with $\delta=10^{-5}$ and the non-private ($\epsilon=\infty$) cases. The models are trained using the DP-SGD algorithm (Abadi et al., 2016) using Opacus (Yousefpour et al., 2021), which we tune for the highest train accuracy on learning rate $lr$, number of epochs $e$, batch size $bs$ and maximum $\ell_{2}$ clipping norm (C) for the largest final accuracy.
For the text data modality, we evaluate the performance of PANORAMIA on the differentially private GPT-2-Small model, which has been fine-tuned on the WikiText dataset (the preprocessing steps are identical to those in Section C.2). As with the image modality, we use various target privacy budgets ($\epsilon$) with $\delta=10^{-5}$, as well as non-private settings ($\epsilon=\infty$). We train the models using the DP-SGD algorithm Abadi et al. (2016), incorporating Ghost clipping Li et al. (2021) for calculating the per-example gradient. We minimize the validation loss by tuning hyperparameters such as the learning rate ($lr$), number of epochs ($e$), batch size ($bs$), and clipping norm ($C$). The lowest validation loss values obtained are $3.79$, $3.7$, and $3.65$ for $\epsilon$ values of 1, 3, and 10, respectively. For the non-private GPT-2-Small model, the lowest validation loss achieved is $3.341$.
The noise multiplier $\sigma$ is computed given $\epsilon$, number of epochs and batch size. We use both PANORAMIA and O(1) (Steinke et al., 2023) privacy loss measurements under the pure $\epsilon$-DP analysis.
@@
Results. For the image data modality, tables 7 and 8 summarizes the auditing results of PANORAMIA on different DP models. For ResNet-18, we observe that at $\epsilon=1,2,4,6$ (more private models) PANORAMIA detects no privacy loss, whereas at higher values of $\epsilon=10,15,20$ (less private models) and $\epsilon=\infty$ (a non-private model) PANORAMIA detects an increasing level of privacy loss with $\tilde{\epsilon}_{\epsilon=10}<\tilde{\epsilon}_{\epsilon=15}<\tilde{\epsilon}_{\epsilon=20}<\tilde{\epsilon}_{\epsilon=\infty}$, suggesting a higher value of $\epsilon$ correspond to higher $\tilde{\epsilon}$. We observe a similar pattern with Wide-ResNet-16-4, in which no privacy loss is detected at $\epsilon=1,2$ and higher privacy loss is detected at $\epsilon=10,15,20,\infty$. We also compare the auditing performance of PANORAMIA with that of O(1) (Steinke et al., 2023), with the conclusion drawn by these two methods being comparable. For both ResNet-18 and Wide-ResNet-16-4, O(1) reports values close to 0 (almost a random guess between members and non-members) for $\epsilon<10$ DP models, and higher values for $\epsilon=10,15,20,\infty$ DP models. The results suggest that PANORAMIA is potentially an effective auditing tool for DP models that has comparable performance with O(1) and can generalize to different model structures.
For the text data modality, Table 9 summarizes the auditing results of PANORAMIA on differentially private (fine-tuned) Large Language Models. In all cases, neither PANORAMIA not O(1) Steinke et al. (2023) do not detect any privacy loss.
@@
ResNet18 $\epsilon=\infty$ PANORAMIA RM;GN 2.44 3.87 1.43 -
O (1) RM;RN - - - 1.471
ResNet18 $\epsilon=20$ PANORAMIA RM;GN 2.44 3.6331 1.19 -
@@
ResNet18 $\epsilon=15$ PANORAMIA RM;GN 2.44 3.57 1.13 -
@@
ResNet18 $\epsilon=10$ PANORAMIA RM;GN 2.44 2.70 0.26 -
O (1) RM;RN - - - 0.28
ResNet18 $\epsilon=6$ PANORAMIA RM;GN 2.44 2.478 0.038 -
@@
ResNet18 $\epsilon=4$ PANORAMIA RM;GN 2.44 1.99 0 -
@@
ResNet18 $\epsilon=2$ PANORAMIA RM;GN 2.44 1.709 0 -
O (1) RM;RN - - - 0.05
ResNet18 $\epsilon=1$ PANORAMIA RM;GN 2.44 1.412 0 -
@@
Target model Audit $\mathbf{c_{\textnormal{lb}}}$$\mathbf{\{\varepsilon+c\}_{\textnormal{lb}}}$ $\mathbf{\tilde{\varepsilon}}$ $\mathbf{\varepsilon}$
WRN-16-4 $\epsilon=\infty$ PANORAMIA RM;GN 2.44 3.02 0.58 -
@@
Target model $f$ Audit Generator $G$ $\mathbf{\tilde{\epsilon}}$ $\mathbf{\epsilon}$
$\infty$-DP GPT2-Small (GPT2_E37) PANORAMIA GPT2-Small $0.22\pm 0.37$
O (1) RM;RN - - $2.82\pm 0.31$
$1$-DP GPT2-Small PANORAMIA GPT-2-Small 00 -
O (1) RM;RN - - 0
$3$-DP GPT2-Small PANORAMIA GPT-2-Small 00 -
O (1) RM;RN - - 0
$10$-DP GPT2-Small PANORAMIA GPT-2-Small 00 -
O (1) RM;RN - - 0
Table 9: Privacy audits on GPT-2-Small target models with varying privacy guarantees.